OpenStudy (anonymous):

LOG[27](9)=x solve for x

OpenStudy (campbell_st):

use some index notation to write 9 in terms of 27 so \[9 = \sqrt[3]{27^2}\] simplify the power for the value of x \[\log_{27} \sqrt[3]{27^2} = x\] hope this helps

OpenStudy (anonymous):

Thank you.